# Helfrich cylinder Alexander Meiners, Hannes Uecker

### v1

@@ 3.2 Center manifold construction
At ${\cal C}_{L}(r)$, $\det{\cal D}=0$, and we cannot perturb ${\cal C}_{L}(r)$ while preserving the area and the volume. Therefore we start with
@@
In §3.3 we will see that there are parameter values $\lambda_{2}$ such that the spectrum of the linearization of (42) is contained in the negative half plane, with one zero eigenvalue with multiplicity according to (44). The spectrum on the finite cylinder is discrete, with a spectral gap between the zero and first non zero eigenvalue. Hence the spectrum decomposes into the center spectrum $\mathrm{spec}_{c}$ and the stable spectrum $\mathrm{spec}_{s}$. Using the well-posedness Theorem 2.2 we may apply [Sim95, Theorem 4.1] on the existence of a center manifold under the above conditions. Since (44) shows that the center eigenspace coincides with the tangent space of the manifold of translated cylinders, both manifolds also coincide locally. This gives the following results.
[…]
At the trivial solution $(0,\lambda_{2})$ with $\lambda_{1}$ and $\lambda_{2}$ related by (2), there exists a smooth center manifold $\mathcal{M}^{c}$ for (42) at least of dimension two. If the kernel of ${\partial}_{(u,\lambda_{2})}{\cal G}_{\cal A}$ is spanned by $\Psi_{0\pm 1}$ then
${\cal M}^{c}=\text{Shifts in $y$ and $z$ direction}.$ (45)
@@
If $\lambda_{2}\in{\mathbb{R}}$ such that the spectrum of ${\partial}_{(u,\lambda_{2})}{\cal G}_{\cal A}$ is contained in $(-\infty,0]$ and $\ker({\partial}_{(u,\lambda_{2})}{\cal G}_{\cal A})$ is spanned by $\Psi_{0\pm 1}$, then there exist $\delta,\omega,M>0$ such that for all $u_{0}\in V^{m}_{\varepsilon}$ with $\|u_{0}\|_{H^{m}}<\delta$ the solution $U=(u,\lambda_{1})$ of (42) exists globally in time and there exist a unique $U_{\infty}\in{\cal M}^{c}$ such that
@@ 3.3 Stability and instability
Following Theorem 3.1, to check for stability of ${\cal C}_{L}(r)$ we seek $\Lambda\in{\mathbb{R}}^{2}$ such that all eigenvalues $\mu(m,n,\Lambda)$ are in $(-\infty,0]$, and the only zero eigenvalues belong to $\Psi_{0\pm 1}$ (translations in $y$ and $z$). From here on, to simplify notation we fix $r=1$. Practically, we vary $\lambda_{2}$ along the trivial branch, and monitor the signs of the eigenvalues. Due to the three parameters $(L,c_{0},\lambda_{2})$, with $\lambda_{1}$ given by (2), we will distinguish between several cases, while the main questions are:
@@
The zeros of $\mu$ give values $\lambda_{2}(m,n)$, indicating which mode induces the instability. Thus we compute the zeros of $\mu(m,n,\Lambda)$ and the slope ${\partial}_{\lambda_{2}}\mu(m,n,\Lambda)$, giving us on which side the eigenvalue is negative, splitting the $(L,c_{0})$ plane. As a visualization, Figure 2 shows plots of the regions of stability (in green) and some spectral plots of selected dispersion relations.
For $m=0$ we have
$\mu(0,n,\Lambda)=-\frac{n^{4}}{2}+\left(\lambda_{2}+1\right)n^{2}-\left(\lambda_{2}+\frac{1}{2}\right),$ (46)
with $\mu(0,\pm 1,\Lambda)=0$ for all $\Lambda$ (translational eigenvalues), and
${\partial}_{\lambda_{2}}\mu(0,n,\Lambda)=n^{2}-1$
is positive for $n>1$. Furthermore $\mu$ is zero at $\lambda_{2}(0,n)=(1-n^{2})^{-1}(-\frac{n^{4}}{2}+n^{2}-\frac{1}{2})$, which is monotonously increasing in $n$. Hence under axially constant perturbations $\mathcal{C}_{r}$ is stable for $\lambda_{2}<\frac{3}{2}$.
@@
Refer to caption Refer to caption
@@
[…]
Refer to caption Refer to caption
Figure 2: (a–c) Instability curves in the $L$$\lambda_{2}$ plane ($\lambda_{1}$ according to (2)) for different values of $c_{0}$, delimiting the $\lambda_{2}$ parameter range for stability of ${\cal C}_{L}$. The blue curve correspond to the mode $(m,n)=(1,0)$ (minimal wave number pearling instability), the orange to $(m,n)=(1,1)$ (minimal wave number coiling/buckling) and the violet $\lambda_{2}=3/2$ to $(m,n)=(0,2)$ (wrinkling, independent of $L$). Lighter blue in (c) correspond to $(m,n)=(2,0)$, which gives the only finite wave number (independent of $L$) instability. (d) illustrates the long wave nature of the (1,0) pearling instability at $(c_{0},L)=(0,30)$ (as $k_{m}=0$ is forbidden by the area (or volume) constraint, and the finite wave number nature of the (2,0) pearling instability at $(c_{0},L)=(0.48,30)$.
Next we consider the modes $m\neq 0\neq n$, in particular $(m,n)=(1,1)$, where
$\mu(1,1,\Lambda)$ $=-\frac{1}{2}((k_{1}^{2}+1)^{2}+(\lambda_{2}+1)(k_{1}^{2}+1)-\left(\lambda_{2}+\frac{1}{2}\right)-(1-2c_{0})k_{1}^{2}$
$=-\frac{1}{2}(k_{1}^{2}+1)^{2}+(\lambda_{2}-2c_{0})k_{1}^{2}+\frac{1}{2}.$
One sees directly that $\mu(1,1,\Lambda)$ is strictly increasing in $\lambda_{2}$ and has a zero for
$\lambda_{2}^{(1,1)}=\frac{1}{2}k_{1}^{2}+1-2c_{0},$ (47)
where $\lambda_{2}$ is strictly decreasing in $L$ because so is $k_{m}=2m\pi/L$. Hence, if $c_{0}>-1/4$, then there exists an $L$ large enough so that the cylinder destabilizes due to the $(m,n)=(1,1)$ mode with $L$ explicitly computable.
For $m\geq 1$ and $n=0$ (pearling), the eigenvalues
$\mu(m,0,\Lambda)=-\frac{1}{2}k_{m}^{4}+(\lambda_{2}+2c_{0})k_{m}^{2}-\lambda_{2}-\frac{1}{2}$
behave as follows. Depending on $L$, either $k_{m}<1$ such that $\mu(m,0,\Lambda)$ is strictly decreasing in $\lambda_{2}$, or $k_{m}>1$ such that $\mu(m,0,\Lambda)$ is strictly increasing in $\lambda_{2}$. In either case, for $m=1$ the eigenvalue has a zero at
$\lambda_{2}^{(m,0)}=\Big{[}1-k_{m}^{2}\Big{]}^{-1}\Bigg{(}-\frac{1}{2}k_{m}^{4}+2c_{0}k_{m}^{2}-\frac{1}{2}\Bigg{)}$ (48)
for all $L\in{\mathbb{R}}\backslash\{2\pi m\}$. First we note that the asymptote at $L=2\pi m$ is odd, where for $c_{0}<\frac{m}{2}$, $\lim_{k_{m}\searrow 1}\lambda_{2}=-\infty$ and for $c_{0}>\frac{m}{2}$ the other way around. In the latter case there always exists a $L$ such that the $m$’th mode destabilizes the cylinder, so we do not focus further on that case. A straightforward analysis of (48) shows that it admits two positive solutions
$L=\begin{cases}L_{G}=2\pi\left(\frac{1+\sqrt{2}\sqrt{1-2c_{0}}}{4c_{0}-1}\right)^{1/2}\quad\text{ for }c_{0}\in(1/4,1/2],\\ L_{S_{1}}=2\pi\left(\frac{1-\sqrt{2}\sqrt{1-2c_{0}}}{4c_{0}-1}\right)\quad\text{ for }c_{0}\in(-\infty,1/2],\\ L_{S_{2}}=\sqrt{2}\pi,{\hskip 56.9055pt}\text{for }c_{0}=1/4,\end{cases}$
where for $i=1,2$ it holds that $L_{S_{i}}<2\pi$ and are minima. For $c_{0}\leq\frac{1}{2}$ and $L>2\pi$ the eigenvalue $\mu(1,0,\Lambda)$ has negative slope at (48), hence for smaller $\lambda_{2}$ the cylinder is destabilized. In the case of $c_{0}\leq\frac{1}{4}$, $\lambda_{2}$ is monotonously increasing in $L$, yielding a maximum wavelength instability. In the regime where $L_{G}$ exists, the ${\cal C}_{L}$ destabilizes to a finite wave length.
In the following Corollary of Theorem 3.1 we summarize the above, where maximal wave length (or long wave instability) refers to the minimal allowed $k_{1}=2\pi/L$, since $k_{0}=0$ is prohibited by the volume and area constraints.
@@
    $c_{0}\in(1/4,1/2)$ and $\lambda_{2}\in(\lambda_{2}^{(1,0)},\lambda_{2}^{(1,1)})$; ${\cal C}_{L}$ destabilizes at $\lambda_{2}^{(1,0)}$ with finite wave length $L_{G}$ wrt the mode $\psi_{20}$ (pearling), and at $\lambda_{2}^{(1,1)}$ wrt the mode $\psi_{11}$ (buckling and coiling) of maximal wave length.
@@
    $c_{0}\in[-1/4,1/4]$ and $\lambda_{2}\in(\lambda_{2}^{(1,0)},\lambda_{2}^{(1,1)})$; ${\cal C}_{L}$ destabilizes at $\lambda_{2}^{(1,0)}$ wrt to the mode $\psi_{10}$ (pearling), and at $\lambda_{2}^{(1,1)}$ wrt the mode $\psi_{11}$ (buckling and coiling), both of maximal wave length.
@@
    $c_{0}\in(-\infty,-1/4]$ and $\lambda_{2}\in(\lambda_{2}^{(1,0)},\lambda_{2}^{(0,2)})$; ${\cal C}_{L}$ destabilizes at $\lambda_{2}^{(1,0)}$ wrt the mode $\psi_{10}$ (pearling) with maximal wave length, and at $\lambda_{2}^{(0,2)}$ wrt the mode $\psi_{01}$ (wrinkling, homogeneous in $x$).
If $c_{0}>1/2$ there is always a positive eigenvalue for every $\lambda_{2}$, hence the cylinder is spectrally unstable.
@@ 4 Bifurcations from ${\cal C}_{L}$
We now compute the branches bifurcating at the instabilities given in Corollary 3.2. In §4.1 we give the general symmetry based bifurcation results, and in §4.2 we derive the associated amplitude equations and moreover discuss the stability of the bifurcating branches.
@@ 4.1 General bifurcation results
00 0 $={\cal A}(u)-{\cal A}_{0}.$ (49b)
The linearization of (49) is the first $2\times 2$ block of ${\partial}_{U}{\cal G}$, hence with eigenvalues (39) with $(m,n)\in{\mathbb{Z}}^{2}\backslash\{(0,0),(0,\pm 1)\}$, and eigenvectors
@@
Here $\mu_{00}$ is not present since it corresponds to changes of radius along the cylinder branch. Solving (2) gives (possible) bifurcation points (for completeness with $r$–dependence)
@@
At (51) in the system (49) we have equivariant pitchfork bifurcations of steady state solutions $X(\varepsilon)=\mathcal{C}_{r}(L)-u(\varepsilon)N$, $\Lambda(\varepsilon)=\Lambda_{0}+{\cal O}(\varepsilon^{2})$, for some $0<|\varepsilon|<\varepsilon_{0}$ sufficiently small. The maps $\varepsilon\mapsto(u(\varepsilon),\Lambda(\varepsilon))\in H^{4}\times{\mathbb{R}}^{2}$, or, equivalently $\varepsilon\mapsto(X(\varepsilon),\Lambda(\varepsilon))\in H^{4}_{\text{per}}(\mathcal{C}_{r}(L),{\mathbb{R}}^{3})\times{\mathbb{R}}^{2}$ are analytic, and we characterize the bifurcations as follows:
@@
    $u=\left(A_{m}e^{{\rm i}k_{m}x}+\mathrm{c.c.}\right)\varepsilon+{\cal O}(\varepsilon^{2})$ (52)
    with some amplitude $A_{m}\in{\mathbb{C}}$, equivariant wrt translations in $x$.
@@
    $u=\left(A_{n}e^{{\rm i}n\varphi}+\mathrm{c.c.}\right)\varepsilon+{\cal O}(\varepsilon^{2})$ (53)
    with some amplitude $A_{n}\in{\mathbb{C}}$, equivariant wrt rotations in $\varphi$.
@@
    $m\neq 0\neq n$ (buckling and coils): Two types branches bifurcate. Buckling
    $u=\left(A_{mn}e^{{\rm i}(k_{m}x+n\varphi)}+\mathrm{c.c.}\right)\varepsilon+{\cal O}(\varepsilon^{2}),$ (54)
    equivariant wrt to translations in $x$ and rotations in $\phi$. The other branch are coils as equal amplitude superpositions of buckling,
    $u=\left(A_{mn}e^{{\rm i}(k_{m}x+n\varphi)}+A_{m-n}e^{{\rm i}(k_{m}x-n\varphi)}+\mathrm{c.c.}\right)\varepsilon+{\cal O}(\varepsilon^{2}),$ (55)
    equivariant wrt to translations in $x$, or, equivalently wrt rotations in $\phi$, which yield the same group orbit.

### v2

@@ 3.2 Center manifold construction
Based on the local existence Theorem 2.6 and the spectral computations from Lemma 3.1 we construct the center manifold for the flow near ${\cal C}_{L}(r)$ which will be the basis for the subsequent stability and bifurcation results. At ${\cal C}_{L}(r)$, $\det{\cal D}=0$, and we cannot perturb ${\cal C}_{L}(r)$ while preserving the area and the volume. Therefore we start with
@@
In §3.3 we will see that there are parameter values $\lambda_{2}$ such that the spectrum of the linearization of (41) is contained in the negative half plane, with one zero eigenvalue with multiplicity according to (43). The spectrum on the finite cylinder is discrete, with a spectral gap between the zero and first non zero eigenvalue. Hence the spectrum decomposes into the center spectrum $\mathrm{spec}_{c}$ and the stable spectrum $\mathrm{spec}_{s}$. Using the well-posedness Theorem 2.6 we may apply [Sim95, Theorem 4.1] on the existence of a center manifold under the above conditions. Since (43) shows that the center eigenspace coincides with the tangent space of the manifold of translated cylinders, both manifolds also coincide locally. This gives the following results, where again $y$ and $z$ are the euclidian coordinates orthogonal to the cylinder axis $x$.
Lemma 3.5.
At the trivial solution $(0,\lambda_{2})$ with $\lambda_{1}$ and $\lambda_{2}$ related by (7), there exists a smooth center manifold $\mathcal{M}^{c}$ for (41) of dimension at least two. If the kernel of ${\partial}_{(u,\lambda_{2})}{\cal G}_{\cal A}$ is spanned by $\Psi_{0\pm 1}$ then
${\cal M}^{c}=\text{Shifts of ${\cal C}_{L}$ in $y$ and $z$ direction}.$ (44)
@@
If $\lambda_{2}\in{\mathbb{R}}$ such that the spectrum of ${\partial}_{(u,\lambda_{2})}{\cal G}_{\cal A}$ is contained in $(-\infty,0]$ and $\ker({\partial}_{(u,\lambda_{2})}{\cal G}_{\cal A})$ is spanned by $\Psi_{0\pm 1}$, then there exists $\delta,\omega,M>0$ such that for all $u_{0}\in V^{m}_{\varepsilon}$ with $\|u_{0}\|_{H^{m}}<\delta$ the solution $U=(u,\lambda_{1})$ of (41) exists globally in time and there exist a unique $U_{\infty}\in{\cal M}^{c}$ such that
@@ 3.3 Stability and instability
Following Theorem 3.6, to check for stability of ${\cal C}_{L}(r)$ we seek $\Lambda\in{\mathbb{R}}^{2}$ such that all eigenvalues $\mu(k_{m},n,\Lambda)$ are in $(-\infty,0]$, and the only zero eigenvalues belong to $\Psi_{0\pm 1}$ (translations in $y$ and $z$). From here on, to simplify notation we fix $r=1$, and recall that the results for general $r$ follow from rescaling $r\mapsto 1$, $L\mapsto rL$ and $c_{0}\mapsto c_{0}/r$. Practically, we vary $\lambda_{2}$, with $\lambda_{1}$ given by (7) along the trivial branch, and monitor the signs of eigenvalues. Depending on $c_{0},L$, the main questions are:
@@
In Lemma 3.1, we fixed L>0L{>}0 yielding the discrete allowed wave numbers $k_{m}=2\pi m/L$, $m\in{\mathbb{Z}}$. Analytically it is more convenient to return to continuous $k\in{\mathbb{R}}$. Thus we consider
$\mu(k,n,\lambda_{2})=-\frac{1}{2}(k^{2}+n^{2})^{2}+(\lambda_{2}+1)(k^{2}+n^{2})-(\lambda_{2}+1)-(1-2c_{0})k^{2},$ (45)
where due to the reflection (in $x$) symmetry of ${\partial}_{u}G(0)$ in (36) only even powers of $k$ appear, and hence we may restrict to $k\geq 0$, and similarly to $n\geq 0$. Subsequently we reintroduce $k_{m}$ to formulate Corollary 3.7. See also Fig.2 for illustration of the results.
(a)                (b)                (c)
Refer to caption Refer to caption Refer to caption
(d)                   (e)
Refer to caption Refer to caption Refer to caption   Refer to caption
@@
For $k=0$ we have
$\mu(0,n,\lambda_{2})=-\frac{n^{4}}{2}+\left(\lambda_{2}+1\right)n^{2}-\left(\lambda_{2}+\frac{1}{2}\right),$ (46)
with $\mu(0,1,\lambda_{2})=0$ for all $\lambda_{2}$ (translational eigenvalues), and ${\partial}_{\lambda_{2}}\mu(0,n,\lambda_{2})=n^{2}-1$ is positive for $n>1$. Furthermore $\mu$ is zero at $\lambda_{2}^{(0,n)}=(1-n^{2})^{-1}(-\frac{n^{4}}{2}+n^{2}-\frac{1}{2})$, which is monotonically increasing in $n$. Hence, under axially constant perturbations ${\cal C}_{L}$ is stable for $\lambda_{2}<\frac{3}{2}$, indicated by the violet line in Fig.2.
Next we consider $k>0$ and $n\neq 0$, in particular $n=1$ which turns out to be the most dangerous $n$ for $k\neq 0$. We have
$\mu(k,1,\lambda_{2})$ $=-\frac{1}{2}((k^{2}+1)^{2}+(\lambda_{2}+1)(k^{2}+1)-\left(\lambda_{2}+\frac{1}{2}\right)-(1-2c_{0})k^{2}$
$=-\frac{1}{2}(k_{1}^{2}+1)^{2}+(\lambda_{2}+2c_{0})k^{2}+\frac{1}{2}=-k^{2}(\frac{1}{2}k^{2}+1-\lambda_{2}-2c_{0}).$
One sees directly that $\mu(k,1,\Lambda)$ is strictly increasing in $\lambda_{2}$ and has a zero at $k>0$ for
$\lambda_{2}=\frac{1}{2}k^{2}+1-2c_{0},$ (47)
indicated by the yellow line in Fig.2. In particular, if $c_{0}>-1/4$, then there exists an explicitly computable $L$ large enough so that the $(k_{1},1)$ coiling/buckling instability, $k_{1}=2\pi/L$, in $\lambda_{2}$ is below the wrinkling instability of ${\cal C}_{L}$.
For $k>0$ and $n=0$ (pearling) we have, with $\kappa:=k^{2}$, i.e., $\tilde{mu}(\kappa,\lambda_{2}):=\mu(k^{2},0,\lambda_{2})$,
$\tilde{mu}(\kappa,\lambda_{2})=-\frac{1}{2}\kappa^{2}+(\lambda_{2}+2c_{0})\kappa-\lambda_{2}-\frac{1}{2},$ (48)
and solving $\mu(\kappa,\lambda_{2})=0$ for $\lambda_{2}=\lambda_{2}(\kappa)$ yields
$\lambda_{2}(\kappa)=\frac{1}{2}\frac{\kappa^{2}-4c_{0}\kappa+1}{\kappa-1}=\frac{1}{2}(\kappa-1)+\frac{(1-2c_{0})\kappa}{\kappa-1}.$ (49)
For $c_{0}\neq\frac{1}{2}$, this has a singularity at $\kappa=1$ with
@@
For $c_{0}=\frac{1}{2}$ we have $\lambda_{2}(\kappa)=-1/2+\kappa/2$. In any case, ${\partial}_{\lambda_{2}}\tilde{mu}(\kappa,\lambda_{2})=\kappa-1$ such that the unstable region is below $\lambda_{2}(\kappa)$ for $0<\kappa<1$ and above $\lambda_{2}(\kappa)$ for $\kappa>1$. See the blue curves/sets in Fig.2. Finally, searching for extrema of $\lambda_{2}(\kappa)$ via ${\partial}_{\kappa}\lambda_{2}(\kappa)=0$ yields
$\kappa_{\pm}=1\pm\sqrt{2-4c_{0}}.$ (50)
Hence, $\kappa_{\pm}$ exist for $c_{0}<1/2$, and $\kappa_{-}\in(0,1)$ for $c_{0}\in(1/4,1/2)$, and (only) in this case we have a finite wave number $k=\sqrt{\kappa}$ instability of ${\cal C}_{L}$, for sufficiently large $L$.
Again, see Fig.2 (a)–(d) for illustration of these instability curves and selected dispersion relations, while in (e) we show the fully unstable case, which will be of interest for the outlook in Fig.14. In the following Corollary we summarize the above, with maximal wavelength refering to the minimal allowed $k_{1}=2\pi/L$.
@@
    $c_{0}\in(1/4,1/2)$ and $\lambda_{2}\in(\lambda_{2}^{\text{pearl}},\lambda_{2}^{\text{cb}})$ where $\lambda_{2}^{\text{pearl}}=\lambda_{2}(\kappa_{-})$ with $\kappa_{-}=1-\sqrt{2-4c_{0}}$ and $\lambda_{2}$ from (49), and where $\lambda_{2}^{\text{cb}}=1-2c_{0}$. In particular, ${\cal C}_{L}$ with large $L$ destabilizes near $\lambda_{2}^{\text{pearl}}$ to a finite wavelength pearling, and near $\lambda_{2}^{\text{cb}}$ to maximal wavelength buckling and coiling.
@@
    $c_{0}\in[-1/4,1/4]$ and $\lambda_{2}\in(\lambda_{2}^{\text{pearl}},\lambda_{2}^{cb})$, where $\lambda_{2}^{\text{pearl}}=-1/2$ and $\lambda_{2}^{\text{cb}}=1-2c_{0}$. In particular, ${\cal C}_{L}$ with large $L$ destabilizes near $\lambda_{2}=-1/2$ to maximal wavelength pearling, and near $\lambda_{2}^{\text{cb}}$ to maximal wavelength coiling and buckling.
@@
    $c_{0}\in(-\infty,-1/4]$ and $\lambda_{2}\in(\lambda_{2}^{\text{pearl}},\lambda_{2}^{\text{wri}})$, where $\lambda_{2}^{\text{pearl}}=-1/2$ and $\lambda_{2}^{\text{wri}})=3/2$. In particular, ${\cal C}_{L}$ with large $L$ destabilizes near $\lambda_{2}=-1/2$ to maximal wavelength pearling, and at $\lambda_{2}=3/2$ to wrinkling (homogeneous in $x$).
If $c_{0}>1/2$, then for all $\lambda_{2}\in{\mathbb{R}}$ there exists an $L_{0}>0$ such that ${\cal C}_{L}$ with $L>L_{0}$ is spectrally unstable.
Remark 3.8.
The values $\lambda_{2}^{\text{cb}}=1-2c_{0}$ and $\lambda_{2}^{\text{pearl}}=-1/2$ are the asymptotics for $k_{1}\to 0$, i.e., for the infinitely long cylinder. In the following section we return to some finite $L>0$, i.e., some finite $k_{1}=2\pi/L$, and in the numerics in §5 we use, e.g., $L=10$, hence $k_{1}=\pi/5\approx 0.628$, or $L=30$, hence $k_{1}=\pi/15\approx 0.209$. The corresponding values for $\lambda_{2}^{\text{cb}}$ and $\lambda_{2}^{\text{pearl}}$ (maximum wavelength pearling) can then be read off from the orange and blue lines in Fig.2. For, e.g., $c_{0}=0.48$ and $L=30$, the finite wavelength pearling corresponds to $k_{m}=2\pi m/30\approx 0.838$ for $m=4$. $\rfloor$
@@ 4 Bifurcations from ${\cal C}_{L}$
We now compute the branches bifurcating at the instabilities given in Corollary 3.7. In §4.1 we give the general bifurcation results, based on the $O(2)\times O(2)$ symmetry of the problem at ${\cal C}_{L}$, and in §4.2 we derive the associated amplitude equations (AEs) and moreover discuss the stability of the bifurcating branches. [^5] We exclude (leave for future work) the exceptional (co–dimension two) cases of combinations $(L,c_{0})$ such that at some $\lambda_{2}$ two different modes become unstable simultaneously. This can for instance happen for large $L$ and $c_{0}\stackrel{{\scriptstyle>}}{{\approx}}-0.25$, namely mode crossing of wrinkling and coiling/buckling.
@@ 4.1 General bifurcation results
$0$ $={\cal A}(u)-{\cal A}_{0}.$ (51b)
[…]
@@
Here $\mu_{00}$ is not present since it corresponds to changes of radius along the cylinder branch. Solving (2) gives (possible) bifurcation points (for completeness with rr–dependence)
@@
At (53) in the system (51) we have equivariant pitchfork bifurcations of steady state solutions $X(\varepsilon)={\cal C}_{L}-\varepsilon u(\varepsilon)N$, $\Lambda(\varepsilon)=\Lambda_{0}+{\cal O}(\varepsilon^{2})$, for $0<|\varepsilon|<\varepsilon_{0}$ sufficiently small. The maps $\varepsilon\mapsto(u(\varepsilon),\Lambda(\varepsilon))\in H^{4}\times{\mathbb{R}}^{2}$, or, equivalently $\varepsilon\mapsto(X(\varepsilon),\Lambda(\varepsilon))\in H^{4}_{\text{per}}({\cal C}_{L},{\mathbb{R}}^{3})\times{\mathbb{R}}^{2}$ are analytic, and we characterize the bifurcations as follows:
@@
    $u=\left(A_{m}e^{{\rm i}k_{m}x}+\mathrm{c.c.}\right)+{\cal O}(\varepsilon)$ (54)
    with amplitude $A_{m}\in{\mathbb{C}}$; the isotropy class can be represented by the isotropy subgroup $\Sigma:=\{\sigma\in\Gamma:\sigma X=X\}=\{(0,m_{x},\tau_{\phi},m_{\phi})\}\leqslant O(2)\times O(2)$, where $m_{x}$ stands for reflections in $x$ at $x_{0}$ with maximal or minimal $r(x_{0})$, $\tau_{\phi}$ stands for translations in $\phi$ (rotations), and $m_{\phi}$ for reflections in $\phi$. The fixed point subspace ${\rm Fix}\ \Sigma$ consists of surfaces of revolution.
@@
    $u=\left(A_{n}e^{{\rm i}n\varphi}+\mathrm{c.c.}\right)+{\cal O}(\varepsilon)$ (55)
    with amplitude $A_{n}\in{\mathbb{C}}$; an isotropy subgroup is $\Sigma=\{(\tau_{x},m_{x},0,m_{\phi})\}$ (translations and reflections in $x$, and reflections in $\phi$). ${\rm Fix}\ \Sigma$ consists of $x$-independent surfaces.
@@
    $m\neq 0\neq n$ (coiling and buckling): Two types of branches bifurcate. Coiling,
    $u=\left(A_{mn}e^{{\rm i}(k_{m}x+n\varphi)}+\mathrm{c.c.}\right)+{\cal O}(\varepsilon),$ (56)
    with $\Sigma=Z_{2}$ generated by joint reflections in $x$ and $\varphi$ (at suitable $x_{0},\varphi_{0})$, and continuous symmetry $\tau_{\phi,-k_{m}\xi/n}\circ\tau_{x,\xi}$, i.e., spatial translation followed by angular rotation. In particular, spatial translation and angular rotation yield the same non–trivial group orbits.
    The other branch corresponds to buckling as equal amplitude superpositions of coiling,
    $u=\left(A_{mn}e^{{\rm i}(k_{m}x+n\varphi)}+A_{m-n}e^{{\rm i}(k_{m}x-n\varphi)}+\mathrm{c.c.}\right)+{\cal O}(\varepsilon),$ (57)
    with $\Sigma=Z_{2}\times Z_{2}$ (independent reflections in $x$ or $\varphi$ at suitable $x_{0},\varphi_{0}$), and no continous symmetries left.